A 4.626 g sample of a hydrocarbon, upon combustion in a combustion analysis apparatus, yielded 6.484 g of carbon dioxide. Calculate the mass percent of carbon in the hydrocarbon.

1
Expert's answer
2022-03-16T09:01:03-0400

"n(CO_2)=\dfrac{mass}{molar\ mass}=\dfrac{6.484\ g}{12 +2*16\ g/mol}=0.1474\ mol"


All of the carbon from the hydrocarbon is converted to carbon dioxide:


"n(C)=n(CO_2)=0.1474\ mol"

"m(C) =n(C)*Ar(C)=0.1474\ mol*12\ g/mol=1.768\ g"


Mass percent of carbon in the hydrocarbon:


"\omega(C)=\dfrac{m(C)}{m(hydrocarbon)}*100\% =\dfrac{1.768\ g}{4.626\ g}*100\%=38.2\%"